OpenStudy (anonymous):

find the vector v that has a magnitude of 4 and is in the same direction of u, where u=<5,-1>

OpenStudy (loser66):

@e.mccormick

OpenStudy (e.mccormick):

Have you done anything with normalizing vercors to get the unit length?

OpenStudy (loser66):

@e.mccormick what do you get? the answer?

OpenStudy (loser66):

take time, friend,

OpenStudy (e.mccormick):

\(4\frac{\vec{u}}{\|\vec{u}\|}\) by finding the normalized version of the vector, you have the same direction and in a unit vector. Then multiply that by 4 units.

OpenStudy (loser66):

yes, I am with you.

OpenStudy (e.mccormick):

100? \[\frac{4}{\sqrt{5^2+(-1)^2}}(5,-1) \implies \left(\frac{20}{\sqrt{26}}, -\frac{4}{\sqrt{26}} \right) \]

OpenStudy (e.mccormick):

And don't forget the - there.

OpenStudy (loser66):

ok, you are right, my bad.

OpenStudy (anonymous):

thanks you guys but i got the answer! lololol.

OpenStudy (e.mccormick):

Hehe. We were just discussing it after the fact. =) And as long as you know the process, you should be able to solve any of them.

OpenStudy (loser66):

@BeautyQueen327 glad to hear it, but it doesn't relate to that any more, we discuss to study, too.

OpenStudy (anonymous):

ok ok sorry